OpenStudy (anonymous):

simplify 2(6x-5)^2+2

OpenStudy (anonymous):

What does \[(6x-5)^2\] mean?

OpenStudy (anonymous):

36x-25

OpenStudy (anonymous):

Sure?

OpenStudy (anonymous):

Remember when it's squared, cubed, etc, it means that there's "that many" of the certain expression. \[x^2 = x*x\] same thing applies for (6x-5)^2 \[(6x-5)^2 = (6x-5)(6x-5)\]

OpenStudy (anonymous):

o ok i forgot.. thanks

OpenStudy (anonymous):

Np